A certain doctor's age is now seven-ninth's of what it will be 18 years from now. Just how old is he now?
Answer by macston(5194)

a=his age
a=7/9(a+18)
a=(7/9)a+14 Subtract (7/9)a from each side.
(2/9)a=14 Multiply each side by 9/2.
a=63 ANSWER: he is currently 63 years old
CHECK:
63=(7/9)(63+18)
63=(7/9)81
63=63